OpenStudy (anonymous):

Which shows 51^2 - 49^2 being evaluated using the difference of squares method? 51^2 - 49^2 = 2601 - 2401 = 200 51^2 - 49^2 = (51 + 49)(51 - 49) = (100)(2) = 200 51^2 - 49^2 = (51 - 49)^2 = 2^2 = 4 51^2 - 49^2 = (2601 + 2401)(2601 - 2401) = 1,000,400

OpenStudy (aum):

\(\large a^2 - b^2 = (a+b)(a-b)\) a = 51, b = 49.

OpenStudy (anonymous):

Okay?

OpenStudy (haseeb96):

A IS THE CORRECT ANSWER

OpenStudy (anonymous):

Thanks

OpenStudy (aum):

\(\large a^2 - b^2 = (a+b)(a-b)\) Therefore, \(\large 51^2 - 49^2 = (51+49)(51-49) = 100*2 = 200\) The correct answer is B.